how many milliliters of a 0.25 M HCl solution would be used to prepare 100.0 mL of a 0.050 M solution of the acid?
Answer #1

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 0.25 M

M2 = 0.05 M

V2 = 100 mL

use:

M1*V1 = M2*V2

V1 = (M2 * V2) / M1

V1 = (0.05*100)/0.25

V1 = 20 mL

Answer: 20 mL